A faulty thermometer has its fixed point 5 and 95 .this thermometer reads the temperature of a body as 59 .then the correct temperature of the body on the Celsius scale is

Grade:12

2 Answers

U.Divyatejasree
7 years ago
Here faulty reading (x)=59
Lower fixed point of faulty readind (L.F.P)=5
Upper fixed point of faulty reading (U.F.P.)=95
Fundamental interval (F.I.)=95-5=90
Correct reading(C)=?
c/100*(x-L.F.P)/F.I.
C/100=54/90=0.6
C=0.6*100=60 celsius scale
Yash Chourasiya
3 years ago
Dear Student

IfTC is the correct temperature on celsisu scale, then
TC−0/100 = (x – lower fixed point)/(upper fixed point - lower fixxed point)
TC/100 = 59−5/95−5
TC/100 = 54/90 = 3/ 5
TC = 300/5 = 60

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya